A spherical straight has a radius of 14.0 cm and a charge of 26.0 uC.Calculate the electric field and the electric potential at(a)r=10.0 cm (b)r=20.0 cm, and (c)r=14.0 cm from the centre.
1
Expert's answer
2020-03-25T11:06:26-0400

a) Since the sphere is conducting, the electric field inside it "r=0.1\ m<R" is zero


"E=0\frac{N}{C}"

"V=8.99\cdot 10^9\frac{26\cdot 10^{-6}}{0.14}=1.67\cdot 10^{6}V"

b)


"E=8.99\cdot 10^9\frac{26\cdot 10^{-6}}{0.2^2}=5.84\cdot 10^{6}\frac{N}{C}"

"V=8.99\cdot 10^9\frac{26\cdot 10^{-6}}{0.2}=1.17\cdot 10^{7}V"

c)


"V=8.99\cdot 10^9\frac{26\cdot 10^{-6}}{0.14}=1.67\cdot 10^{6}V"

"E=8.99\cdot 10^9\frac{26\cdot 10^{-6}}{0.14^2}=1.19\cdot 10^{7}\frac{N}{C}"